OpenStudy (anonymous):

Establish identity 1-cosɵ/1+cosɵ=(cscɵ-cotɵ)²

OpenStudy (experimentx):

\[\frac{1-cosɵ}{1+cosɵ}* \frac{1-cos\theta}{1-cos\theta}\]

OpenStudy (experimentx):

rest should easy

OpenStudy (shayaan_mustafa):

ok tell me which side you want to take to prove this identity? tell me your likeness. then we will follow the steps.

OpenStudy (shayaan_mustafa):

experimentX has taken LHS we can prove it by using RHS good luck.

OpenStudy (experimentx):

i really recommend that you should give your best try ...

OpenStudy (anonymous):

Yeah I just needed some help getting started on this one. Thank you!

OpenStudy (shayaan_mustafa):

you are welcome.

OpenStudy (anonymous):

I am still confused about what to do next

OpenStudy (experimentx):

\[\frac{(a-b)^2}{(a-b)(a+b)}\]

OpenStudy (experimentx):

\[\frac{1^2 - 2 \cos\theta+\cos^2\theta}{1 - \cos^2\theta}\]

OpenStudy (experimentx):

\[\frac{1 - 2 \cos\theta + \cos^2\theta}{\sin^2\theta}\]

OpenStudy (experimentx):

now separate \(\sin^2\theta \) from the denominator, you will get your required result.

OpenStudy (anonymous):

Okay thanks!